# Ray-Singer Torsion, Topological Strings and Black Holes

### 23 Jan 2024 v1

@@ 2 Ray-Singer Torsion in Compact Kahler and Calabi-Yau Manifolds
This talk is based on old and new works. In particular the old work with Bershadsky, Cecotti and Ooguri [9] and the new work with van de Heisteeg, Wiesner and Wu [3, 4]. Road-map for the rest of this note is as follows. In section 3 we explain the sense in which the Planck scale is viewed as the short distance cutoff of quantum gravity theories. In section 4 we briefly explain the Bekenstein-Hawking entropy of black holes. In section 5 we explain why having many light species raises a puzzle and a resolution is offered by replacing the Planck scale cutoff with the Species scale cutoff. In section 6 some aspects of string dualities are reviewed with particular attention to how the light states appear as we approach boundaries of moduli spaces. In section 7 we review the relevance of one-loop topological strings (genus 1 Gromov-Witten invariants) to certain $R^{2}$ curvature corrections in the action. Moreover it is shown how mirror symmetry can be used to relate this to a holomoprhic version of Ray-Singer torsion and how this leads to its computation. We use this to find the dependence of the species scale on moduli which leads to interesting bounds on the allowed potential in physical theories. We also explain why bounds on the shape of potential is an important topic for theoretical physics today. In section 8 we end with some concluding remarks.
@@ 4 Basics of Black Hole Entropy
The entropy signifies the number of micro-states of a black hole. If we assume a black hole is made of $N$ independent constituents then this is the same as expecting $N$ to grow with the radius of the black hole by
@@ 5 Light Species Puzzle for Black Holes
Suppose we have $N\gg 1$ of light ($m\ll M_{pl}$) states. Then for any black hole bigger than Planck scale (so that we can ignore the corrections to the Einstein-Hilbert action), we would expect
$S>N,$
@@ 7 Topological Strings on CY 3-folds and dependence of Species scale on moduli
So to complete the circle of ideas we need to know how to compute $F_{1}$, the genus one topological string amplitude? Here we can use mirror symmetry between Type IIA strings on CY3 with Type IIB on Mirror CY3’ (See [11] for more detail on this). In particular note that
@@
This maps the Kahler parameters of CY $k$ (parameterized by $h^{1,1}$) to complex moduli $t$ of CY’ (parameterized by $h^{1,2}$). Therefore Topological String on CY which is parameterized by Kahler deformations gets mapped to complex deformation of CY’.

### 10 Mar 2024 v2

@@ 2 Ray-Singer Torsion in Compact Kahler and Calabi-Yau Manifolds
This talk is based on old and new works. In particular the old work with Bershadsky, Cecotti and Ooguri [9] and the new work with van de Heisteeg, Wiesner and Wu [3, vandeHeisteeg:2023ubh]. Road-map for the rest of this note is as follows. In section 3 we explain the sense in which the Planck scale is viewed as the short distance cutoff of quantum gravity theories. In section 4 we briefly explain the Bekenstein-Hawking entropy of black holes. In section 5 we explain why having many light species raises a puzzle and a resolution is offered by replacing the Planck scale cutoff with the Species scale cutoff. In section 6 some aspects of string dualities are reviewed with particular attention to how the light states appear as we approach boundaries of moduli spaces. In section 7 we review the relevance of one-loop topological strings (genus 1 Gromov-Witten invariants) to certain $R^{2}$ curvature corrections in the action. Moreover it is shown how mirror symmetry can be used to relate this to a holomoprhic version of Ray-Singer torsion and how this leads to its computation. We use this to find the dependence of the species scale on moduli which leads to interesting bounds on the allowed potential in physical theories. We also explain why bounds on the shape of potential is an important topic for theoretical physics today. In section 8 we end with some concluding remarks.
@@ 4 Basics of Black Hole Entropy
The entropy signifies the number of micro-states of a black hole. If we assume a black hole is made of $N$ independent constituents (so that the total number of states goes as $e^{N}$) then this is the same as expecting $N$ to grow with the radius of the black hole by
@@ 5 Light Species Puzzle for Black Holes
Suppose we have $N\gg 1$ of light ($m\ll M_{pl}$) states. Then for any black hole bigger than Planck scale (so that we can ignore the corrections to the Einstein-Hilbert action), we would expect the entropy of black hole $S_{BH}$ satisfies
$S_{BH}>N,$
@@ 7 Topological Strings on CY 3-folds and dependence of Species scale on moduli
So to complete the circle of ideas we need to know how to compute $F_{1}$, the genus one topological string amplitude? Here we can use mirror symmetry between Type IIA strings on $CY3$ with Type IIB on Mirror $CY3’$ ’ (See [11] for more detail on this). In particular note that
@@
This maps the Kahler parameters of $CY$ $k$ (parameterized by $h^{1,1}$) to complex moduli $t$ of $CY^{\prime}$ (parameterized by $h^{1,2}$). Therefore Topological String on $CY$ which is parameterized by Kahler deformations gets mapped to complex deformation of $CY^{\prime}$.